### v1

# Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer

Abstract
We propose Rotate, Clip, and Partition (RCP), a Quantization-Aware Training (QAT) approach that first realizes extreme compression of LLMs with W2A4KV4(2-bit weight, 4-bit activation, and 4-bit KV-cache) configuration. RCP integrates recent rotation techniques with a novel non-uniform weight quantizer design, by quantitatively analyzing the impact of random rotation on 2-bit weight quantization. Our weight quantizer features Learnable Direct Partitioning (LDP), which introduces learnable parameters to directly learn non-uniform intervals jointly with LLM weights. We also present a specialized GPU kernel that supports GEMV on non-uniform W2A4. Experiments show that RCP can compress LLaMA-2-7B to W2A4KV4 with a loss of only 2.84 WikiText2 ppl and 5.29 times reduced memory footprint. Furthermore, RCP can quantize challenging mobile-targeted LLaMA-3.2 models and domain-specific WizardCoder-7B and MetaMath-7B with no critical problems such as convergence failure and repetition. Code will be made available at blind_review.
[…]
3pt
Rotate, Clip, and Partition: Towards W2A4KV4 Quantization by Integrating Rotation and Learnable Non-uniform Quantizer
Euntae Choithanks: these authors contributed equally., Sumin Song11footnotemark: 1, Woosang Lim, Sungjoo Yoo Seoul National University euntae.choi175@gmail.com, songsm921@snu.ac.kr, ftyg656512@snu.ac.kr, sungjoo.yoo@gmail.com
1 Introduction
Large language models (LLMs) have made significant advancements, but their growing size and resource demands create challenges for deployment across data centers and mobile devices. To address these constraints, extensive research efforts have focused on improving the efficiency of LLM serving through various strategies. Quantization, one of the various methods, has emerged as a particularly straightforward and effective method for reducing memory consumption and inference costs without severely compromising accuracy. By lowering the numerical precision of parameters and intermediate representations, quantization leverages hardware capabilities such as half-precision and integer tensor cores.
[…]
Figure 1: Bit-Level scaling laws for perplexity for LLaMA-3 AI@Meta (2024) (1B, 3B, 8B).
To address these issues, rotation techniques Ashkboos et al. (2024b); Liu et al. (2024b) have proven effective at W4A4KV4. Nevertheless, extending these methods to even lower bit widths, such as W2 or W3, remains insufficient, as the severe information bottleneck cannot be overcome with a uniform quantization and PTQ, ultimately resulting in a loss of generation quality.
Figure 2: A chain of thought (CoT) reasoning example from GSM8K Cobbe et al. (2021) benchmark, conducted by two differently quantized MetaMath-7B Yu et al. (2023) models. The result on the left is from a state-of-the-art QAT method BitDistiller Du et al. (2024). On the right, our proposed RCP is applied. Both methods employ exactly the same 4-bit quantization setting for activation and KV-cache.
In this work, we propose a Quantization-Aware Training (QAT) approach called Rotate, Clip, and Partition (RCP), which integrates rotation techniques with non-uniform quantization strategies to achieve extreme compression of LLMs, specifically realizing W2A4KV4 configurations. Our method jointly optimizes quantization parameters and model weights by systematically incorporating rotation and non-uniform group quantization. We recover meaningful model performance and demonstrate that even 2-bit weight precision, once deemed unattainable, is now within reach. We further extend our techniques to more challenging models, such as LLaMA-3.2 AI@Meta (2024), as well as smaller, mobile-targeted architectures, thereby broadening the applicability of these methods. To facilitate practical deployment, we design efficient W2A4 inference kernels for decoding phase on modern GPUs, thus enabling real-time, resource-efficient LLM serving at extreme compression levels.
Our contributions are summarized as follows:
  • We propose a QAT algorithm that successfully leverages the benefit of random rotation to first achieve W2A4KV4 quantization of LLMs.
  • We demonstrate the scalability of our approach to more challenging (e.g., LLaMA-3.2) and smaller, mobile-targeted models.
  • Additionally, we develop efficient W2A4 inference kernels for the decoding stage on modern GPUs, enabling real-time, resource-efficient LLM serving under extremely low-bit precisions.
2 Preliminaries
2.1 Random Rotation for LLM Quantization
QuaRot Ashkboos et al. (2024b) proposed to apply random rotations while keeping the computational invariance suggested in SliceGPT Ashkboos et al. (2024a). Random rotation suppresses activation outliers and helps quantization, successfully achieving W4A4KV4 with minimal performance loss.
As shown in Fig. 4, $R_{1}$ rotates all transformer decoder layers’ input and output activations and its inverse ($R_{1}^{T}$) is usually fused into the neighboring model weights. $R_{2}$ and $R_{4}$ both require online rotation during inference as they target the intermediate activations in the MHA and FFN layers, respectively. $R_{2}$ is decomposed into two orthogonal matrices: one with head dimension size ($R_{H}$), applied during the V projection, and another with the number of heads ($R_{H}^{\prime}$), applied to the self-attention activation. Its transpose ($R_{2}^{T}$) is applied as a whole to the out-projection in the self-attention layers. $R_{3}$ rotates query (Q) and key (K) vectors after RoPE, ensuring that the KV-cache can be compressed without affecting the self-attention output.
2.2 Asymmetric Quantization and Clipping
Asymmetric quantization uses min/max of data to match data’s asymmetric distribution. This approach can offer higher quantization accuracy than symmetric quantization. The asymmetric quantization equation can be formulated as follows:
$\begin{split}\mathbf{W_{q}}=\mathrm{clamp}(\lfloor\frac{\mathbf{W}}{h}\rceil+z,0,2^{N}-1),\\ \mathrm{where}\,\,h=\frac{\max(\mathbf{W})-\min(\mathbf{W})}{2^{N}-1},\\ z=-\lfloor\frac{\min(\mathbf{W})}{h}\rceil\end{split}$ (1)
Clipping is an optimization technique that truncates outlier values to a predetermined range before quantization. It improves quantization precision within the compressed range, preserving performance while increasing compression ratios. Our proposed method, RCP is based on the learnable weight clipping (LWC) Shao et al. (2024), an asymmetric quantization to obtain asymmetric clipping values that minimize quantization error for each linear layer and then uses them as initial values for learnable clipping $f$ as follows:
$f(\textbf{W},\beta,\gamma)=clip(\textbf{W},\sigma(\beta)\text{min}(\textbf{W}),\sigma(\gamma)\text{max}(\textbf{W}))$ (2)
where $\beta$ and $\gamma$ are learnable parameters for clipping and $\sigma$ is the sigmoid function.
3 Observation: Rotation Sharpens Weight Distribution
(a) Per-group kurtosis measured on the 15th key projection weight in LLaMA-2 7B model before and after random Hadamard transform with layernorm fusion.
(b) Per-group kurtosis measured on the input activation of the 15th key projection weight in the same manner as 3(a) in LLaMA-2 7B model. Due to a large difference of average value, the y axis is presented in log scale. Since almost all groups had small negative kurtosis value after transform, we take the absolute value of them.
Figure 3: Kurtosis analysis results to demonstrate the effect of random rotation on model weight and activation.
As shown in Fig. 1, simply applying rotation techniques before LLM quantization can’t guarantee the performance in extremely low-bit scenarios like W2A4. Building on SmoothQuant Xiao et al. (2023)’s insights, outlier suppression methods improve quantization possibility in the activation space while accepting modest weight space compromises. But recent rotation based PTQ techniques predominantly emphasize the advantages in the activation space, not the disadvantages in the weight space. While rotation-based PTQ techniques work well for moderate quantization up to W4, at W2, the quantization possibility significantly drops.
In this section, we analyze quantization possibility after random Hadamard transformation in both weight and activation spaces in LLaMA-2 7B model. Kurtosis, which measures the fourth standardized moment of a distribution, serves as a key indicator for evaluating outlier frequency and concentration Lee et al. (2024) that intensify the difficulties in quantization. We selected the 15th key projection module for analysis because it showed one of the highest mean squared errors (MSE) when comparing the distributions before and after transformation among all modules. We analyze kurtosis of weight and input activation in group size of 128 following their channels to fully consider the effects of groupwise quantization configuration used in most quantization methods.
Our analysis in Fig. 3(b) reveals that rotation effectively distributes activation values along their channel direction, similar to methods like SliceGPT Ashkboos et al. (2024a), QuaRot Ashkboos et al. (2024b) and PrefixQuant Chen et al. (2024a). The original space shows infinite kurtosis due to extreme outliers, while the rotated space shows predominantly near-zero values. From the perspective of per-channel uniform quantization in the activation, the transformation reduces quantization error compared to the original space.
Conversely, our analysis in Fig. 3(a) shows contrasting circumstances. After going through essential processes such as LayerNorm Fusion and Random Hadamard Transformation, an initially platykurtic distribution with mean kurtosis near 1 becomes more leptokurtic with mean kurtosis near 3.75. These observations indicate that the weight distribution deviates further from the uniform distribution, leading to significant quantization errors when using W2 uniform quantizers, as most weights cluster in two center quantization bins. This finding motivates the development of non-uniform quantizers to optimize bin width distributions. This analysis also explains why FP3 variants outperform INT3 quantization in W3 regime, as NF3’s approach better handles bell-shaped weight distributions, as demonstrated in AFPQ Zhang et al. (2023).
4 Methodology
We propose a non-linear weight quantization (Section 4.1), an efficient dequantization for non-linear weight partitions (Section 4.2), and a GPU inference kernel for extreme low-bit weights (Section 4.4).
Figure 4: An overview of our QAT-based KD framework of RCP. RCP performs rotation-aware clipping, followed by learnable non-uniform quantization (LDP). During this process, knowledge distillation is employed as the training method.
4.1 Differentiable Non-uniform INT2 Group Quantizer
Rotation-aware Clipping Initialization
We initialize clipping parameters in a rotation-aware manner as follows:
$\operatorname*{argmin}_{\beta,\gamma}\|Q(\textbf{W}_{R})\textbf{X}_{R}-\textbf{W}_{R}\textbf{X}_{R}\|^{2}\vspace{-0.2cm}$ (3)
where $\textbf{W}_{R}=R^{T}_{front}f(\textbf{W},\beta,\gamma)R_{rear}$ is the weight matrix obtained by applying appropriate rotations after the clipping $f$ in Eqn. 2. For instance, $R_{front}$ and $R_{rear}$ are set to $R_{1}$ and $I$ for up and gate projections, respectively. Rotation configuration for the other types of weights can be found on the left side of Fig. 4.
Learnable Direct Partitioning
Figure 5: A diagram of Learnable Direct Partitioning (LDP).
The key component of our proposed RCP is a weight quantizer module named learnable direct partitioning (LDP). In order to realize asymmetric weight quantization, we determine value range by LWC and the partitions within the value range by LDP. LDP starts by normalizing the weight W by a dynamic range $h=\sigma(\gamma)\text{max}(\textbf{W})-\sigma(\beta)\text{min}(\textbf{W})$ determined by the LWC. In LDP, two partitioning variables $s_{1}$ and $s_{2}$ are introduced per weight group [^2] For brevity, we do not use a separate notation for group quantization unless necessary since our method can be applied likewise simply by reshaping a weight so that the last dimension becomes the group size., which split the range $h$ into three subsections.
$\begin{split}p_{1}=\sigma(s_{1}),p_{2}=(1-p_{1})\sigma(s_{2}),\\ p_{3}=(1-p_{1})(1-p_{2})\end{split}$ (4)
As shown in Fig. 5 and Eqn. 4, $s_{1}$ defines the portion of the first partition $p_{1}$ in the whole dynamic range $h$. $s_{2}$ defines the portion of the second partition $p_{2}$ from the remaining range $(1-p_{1})h$. The third one is trivially calculated. This formulation guarantees that i) the dynamic range $h$ is seamlessly filled out, ii) each portion is constrained between 0 and 100% via the sigmoid re-parametrization, and iii) no matter how the parameters are updated, the order of the partitions stays the same. The partitioning parameters are initialized as $s_{1}=\sigma^{-1}(1/3)$ and $s_{2}=\sigma^{-1}(1/2)=0$ to evenly split the quantization range in the beginning. The AWQ-styled grid search Lin et al. (2024a) can also be used to find the optimal partition widths; however, the computational burden will grow exponentially since we have to iterate over a four-dimensional loop (two for LWC and two for LDP).
The quantization process of our LDP can then be derived as follows:
$\textbf{W}_{q}=\text{u}(\frac{\textbf{W}}{h}-t_{1})+\text{u}(\frac{\textbf{W}}{h}-t_{2})+\text{u}(\frac{\textbf{W}}{h}-t_{3})$ (5)
where $u(x)$ is the step function and transition point $t_{i}$ (i.e., the right edge of each quantization bin) is set to the center of each partition, computed as $t_{1}=p_{1}/2$ and $t_{i}=t_{i-1}+(p_{i-1}+p_{i})/2$. The straight-through estimator is applied to all step functions so that every parameter (including LLM weights, clipping, and partitioning parameters) can be updated via backpropagation.
Figure 6: An overview of our GPU GEMV kernel with data path along memory hierarchy, pipelining, and epilogue concisely illustrated. wid is the warp index and the per-thread accumulator is simplified (warp lane dimension is not shown).
4.2 Dequantization for Non-Linear Weight Partitions
Unlike the usual uniform quantization scheme, mapping the quantized weights back to real values is not trivial in NU quantizers as the design space of NU dequantization method is large and the inference pipeline is directly affected. In NU2U Liu et al. (2022), the quantized weight $\textbf{W}_{q}$ is simply dequantized to a uniform grid. Similarly, LLT Wang et al. (2022) trains a learnable lookup table for weight and activation to adjust the quantization bin widths but keeps the dequantization the same as other common uniform quantizers.
This design choice has an obvious advantage: Inference can be done on existing hardware without any modification. However, we propose non-uniform dequantization as described in Eqn. 6, based on our observation that uniform dequantization can lead to performance drop under extremely low bit configurations, especially on smaller models.
$\begin{split}\textbf{W}_{deq}=&\sigma(\beta)\text{min}(\textbf{W})\\ +\frac{h}{3}\biggl{(}&\text{u}(\frac{\textbf{W}}{h}-t_{1})(w_{1}-w_{0})\\ +&\text{u}(\frac{\textbf{W}}{h}-t_{2})(w_{2}-w_{1})\\ +&\text{u}(\frac{\textbf{W}}{h}-t_{3})(w_{3}-w_{2})\biggr{)}\end{split}$ (6)
The procedure is designed upon Eqn. 5 with several modifications. The value range is shifted and scaled from $[0,1]$ to $[\sigma(\beta)\text{min}(\textbf{W}),\sigma(\gamma)\text{max}(\textbf{W})]$. The dequantization grid $w_{i}$ is defined in Eqn. 7; the first and last values are the minimum and maximum values of the normalized weight and the middle values are set to the center of the two consecutive transition points.
$w_{i}=\begin{cases}0,&\text{if }i=0\\ \frac{t_{i}+t_{i+1}}{2},&\text{if }0<i<3\\ 1,&\text{if }i=3\end{cases}$ (7)
The dequantization LUT for a quantization unit (a weight group of size 128 in this paper’s experiments) can be pre-computed without any runtime overhead as follows:
$\hat{\textbf{W}}=\{\hat{\text{W}}_{0},\hat{\text{W}}_{1},\hat{\text{W}}_{2},\hat{\text{W}}_{3}\}$ (8)
where $\hat{\text{W}}_{i}=\sigma(\beta)\text{min}(\textbf{W})+h\cdot w_{j}$.
4.3 An NF3 Variant of LDP
We apply not only 2-bit weight quantization but also 3-bit quantization using the asymmetric NF format of AFPQ Zhang et al. (2023) where separate scale values are computed for the negative and positive weights ($s_{neg}=\text{max}(|\textbf{W}_{neg}|)$, $s_{pos}=\text{max}(\textbf{W}_{pos})$). Although shown effective, such NF3 quantizer can lead to suboptimal performance when the distribution is not zero-centered. Therefore, we make a further improvement by applying the proposed LDP to this situation.
The idea is to employ the same learnable clipping parameters ($\beta$, $\gamma$) to obtain the quantization range $h$ and one partitioning parameter $s_{1}$ to express the learnable center point as $c=\sigma(\beta)\text{min}(\textbf{W})+h\cdot\sigma(s_{1})$. Then, the two scale values are updated as follows:
$\begin{split}s_{neg}=|c-\sigma(\beta)\text{min}(\textbf{W})|,\\ s_{pos}=|\sigma(\gamma)\text{max}(\textbf{W})-c|,\end{split}$ (9)
and the quantization process is derived as follows:
$\textbf{W}_{q}=\begin{cases}\lfloor\frac{\textbf{W}-c}{s_{pos}}\rceil,&\text{if }\textbf{W}>c\\ \lfloor\frac{\textbf{W}-c}{s_{neg}}\rceil,&\text{otherwise.}\end{cases}$ (10)
The dequantization is done simply by multiplying the scales to $\textbf{W}_{q}$ and adding $c$.
4.4 W2A4 Look-up Table (LUT) Inference
Designing the inference pipeline of non-uniform W2A4-quantized models poses a big challenge. First, efficient INT tensor cores cannot be utilized since accumulating the multiplication results in INT quantized space makes it impossible to dequantize the weights back to correct non-uniform real values in the LUT $\hat{\textbf{W}}$. Second, both weights and activations must undergo online dequantization to support dynamic quantization, which adds a large amount of computation overhead.
Therefore, we focus on designing GEMV kernel for LUT decoding predominantly bounded by memory bandwidth, which is ideal for featuring the advantage of extreme W2A4KV4 compression. We report our early exploration on GEMM design in Section A.5.
Kernel Arguments and Block Tiling
[…]
Each thread block consists of 128 threads (4 warps) and we only tile along the output dimension and define the tile size as BH. The reason we do not follow the traditional 2-dimensional tiling is that both the input tokens and weights are stored in row-major format and have sub-byte packing along the column direction, which makes it hard to efficiently use high-bandwidth memory that performs best when reading 128B data consecutively. Also, global loads with small transactions and repeated shared stores complicate the pipeline design for latency hiding and degrade overall performance.
Data Path and Latency Hiding Pipeline
[…]
To make our kernel efficient via latency hiding, we design a pipelining strategy where a thread block handles a half of the output elements (BH/2) and iterates twice. At the beginning, an asynchronous copy of $\hat{\textbf{W}}$ and the first $\textbf{W}_{q}$ chunk (of size BH/2 × C/4) is issued using cp.async instruction (1-1 in Fig. 6). Simultaneously, $\textbf{X}_{q}$ is synchronously loaded from global memory and dequantized to be stored into sX (1-2), overlapping activation dequantization latency with loading the first weight chunk.
Subsequently, while we bring in the second $\textbf{W}_{q}$ chunk using cp.async (2-1), we perform dequantization, inner product, and warp reduce on the first $\textbf{W}_{q}$ chunk at the same time (2-2), thereby hiding the second chunk loading latency with computation of the first chunk. Finally, the computation on the second chunk is performed (3) and the shared output array is reduced once more if necessary.
Additional details (e.g., dequantization implementation, shared output) not mentioned here are provided in Section A.6.
#Bits (W-A-KV) Configuration LLaMA-1 7B LLaMA-2 7B LLaMA-3 8B LLaMA-3.2 1B LLaMA-3.2 3B
Method Rotation LDP MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki MMLU 0-shot Wiki
16-16-16 35.10 68.40 5.68 46.45 61.67 5.47 68.40 72.93 6.10 32.20 58.90 9.74 58.00 65.30 7.82
2-4-16 BitDistiller 25.88 42.56 23.19 26.24 43.36 16.47 23.11 39.46 Inf 25.00 36.82 Inf 24.41 37.89 Inf
Ours 26.75 52.28 8.79 26.04 51.49 8.93 29.80 50.59 13.68 25.00 41.08 31.32 29.60 45.29 18.79
Ours 26.98 52.46 8.28 28.04 51.10 8.18 31.87 50.86 12.48 26.30 41.35 27.46 31.40 45.71 16.96
2-4-4 BitDistiller 24.45 43.08 19.98 26.59 44.93 17.40 23.29 39.75 Inf 24.66 37.55 Inf 24.62 37.26 Inf
Ours 26.98 52.21 8.92 26.41 51.10 8.93 29.66 49.80 14.05 24.74 40.77 33.86 31.44 44.26 19.58
Ours 27.34 52.29 8.28 26.92 51.22 8.31 31.01 50.41 12.69 25.62 41.80 29.30 30.33 45.56 17.52
3-4-16 BitDistiller 26.88 55.68 7.47 31.72 56.15 7.04 42.24 55.39 10.19 26.06 37.53 Inf 25.22 37.32 Inf
Ours 28.70 58.52 6.44 34.30 59.28 6.25 54.16 61.06 7.92 26.45 47.88 13.75 47.34 55.66 9.82
Ours 29.46 59.39 6.39 37.33 59.74 6.23 55.33 61.53 7.80 27.77 48.18 13.68 47.31 55.87 9.74
3-4-4 BitDistiller 27.04 56.05 7.54 30.19 55.51 7.15 40.70 56.35 10.46 25.48 38.75 Inf 25.91 37.27 Inf
Ours 28.80 58.48 6.45 33.46 58.53 6.36 51.74 59.69 8.04 26.11 47.14 14.58 46.08 55.08 10.05
Ours 30.00 58.55 6.39 36.07 59.27 6.33 52.55 61.11 7.95 26.54 47.71 14.44 46.40 55.12 9.99
Table 1: Comparison of the perplexity score on WikiText2, MMLU (5s), and 0-shot. 0-shot is average score of 4 Zero-shot Common Sense Reasoning tasks. We show the perplexity results >100 by Inf. Full results of Zero-shot tasks are in the Appendix.
5 Experiments
5.1 Experimental Settings
Models and Tasks
We evaluate RCP on LLaMA-1 Touvron et al. (2023a) 7B, LLaMA-2 Touvron et al. (2023b) 7B, LLaMA-3 AI@Meta (2024)(1B, 3B, 8B). Our evaluation of RCP was carried out on PIQA Bisk et al. (2020), HellaSwag Zellers et al. (2019), WinoGrande Sakaguchi et al. (2021), ARC-c Clark et al. (2018) and MMLU Hendrycks et al. (2020). We use LLM-Humaneval-Benchmarks Chen et al. (2021) and GSM8K Cobbe et al. (2021) for reasoning tasks evaluation. We also report the perplexity score on WikiText2 Merity et al. (2016) for our evaluation.
Training Data
For a fair comparison with our baseline, we use the instruction-tuning data from Alpaca Taori et al. (2023) and the training set of WikiText2 for general language tasks. For understanding and generating code, we use Evol-Instruct-Code [^3] https://github.com/nickrosh/evol-teacher. For math reasoning we use MetaMathQA Yu et al. (2023).
Experiment Configuration
[…]
5.2 Results
Language Modeling Tasks
The results are summarized in Table 1. From the perspective of general language tasks, our method demonstrates the ability to quantize activations and KV-cache under the W2 settings to 4-bit, which was previously unattainable using existing QAT methods. The application of rotation effectively addresses the outlier issues, a common bottleneck in quantization, enabling stable performance even in extremely low-bit quantization scenarios. Furthermore, the addition of LDP not only improves performance on general language tasks but also enhances the accuracy of zero/few shot tasks, which were not adequately addressed by rotation alone. In case of LLaMA-2 7B W2A4KV4, a performance degradation is observed when using rotation only compared to the baseline. However, by incorporating LDP, consistent performance improvements were achieved.
Reasoning Tasks
#Bits (W-A-KV) Configuration WizardCoder 7B MetaMath 7B
Method Rotation LDP HumanEval GSM8K
16-16-16 54.88 66.41
2-4-16 BitDistiller 2.43 0.0
Ours 14.63 1.25
Ours 27.44 41.64
2-4-4 BitDistiller 3.50 5.39
Ours 6.09 0.16
Ours 23.20 40.16
3-4-16 BitDistiller 0.0 0.0
Ours 39.02 0.0
Ours 40.85 54.69
3-4-4 BitDistiller 0.0 0.0
Ours 41.46 0.0
Ours 43.29 52.73
Table 2: Reasoning task results of RCP on domain-specific LLMs.
The results of the reasoning tasks are summarized in Table 2. We evaluate reasoning capabilities in the domains of coding and mathematics.
For the coding domain-specific model, WizardCoder Luo et al. (2023), BitDistiller failed to offer the functional quantized models in both W3 and W2 settings. In our method, applying rotation alone was not effective in W2 settings and recovered some output quality in W3 settings. By incorporating LDP, we achieved up to a threefold improvement in performance, with accuracy increasing from 6.09% to 23.20% under the W2A4KV4 configuration. As shown in Fig. 8 with the application of LDP, we were able to produce logically correct code outputs and eliminate repetition of meaningless code generation.
For the mathematical reasoning model, MetaMath Yu et al. (2023), the baseline BitDistiller and ours without LDP failed to offer functional quantized models while ours with LDP could produce working quantized models. These results highlight the critical role of LDP in enabling proper task performance for reasoning models under extreme low-bit quantization. The output comparison for this task is summarized in Fig. 2.
Inference
Layer Size (2048, 2048) (3072, 3072) (4096, 4096)
FP16 0.042 0.047 0.051
QuaRot 0.077 0.057 0.078
QuaRot+FP16Had 0.158 0.210 0.159
QuaRot+FP32Had 0.194 0.238 0.191
RCP 0.028 0.03 0.040
RCP+FP16Had 0.114 0.167 0.110
RCP+FP32Had 0.136 0.204 0.148
Table 3: GEMV latency without activation quantization overhead. The layer size is composed as (input channel, output channel). All latency numbers are in milliseconds. Full results are in the Appendix.
3.2-1B 3.2-3B 1.2-7B 3-8B
FP16 2.47GB 6.43GB 13.48GB 16.06GB
RCP W3 1.46GB (1.69x) 2.77GB (2.32x) 3.26GB (4.14x) 5.05GB (3.18x)
RCP W2 1.35GB (1.82x) 2.46GB (2.62x) 2.55GB (5.29x) 4.28GB (3.75x)
Table 4: Memory footprint comparison for different weight precisions. Note that 1.2-7B refers to LLaMA-1 and LLaMA-2.
Table 3 and 4 present the results for GEMV in terms of latency and memory comsumption. The latency of GEMV, excluding the activation quantization overhead, is faster compared to FP16 and QuaRot Ashkboos et al. (2024b). This improvement can be attributed to the lower bit precision, which enhances computational efficiency. Table 4 measures the peak memory footprint for W2A4 and W3A4. For W2A4, a significant reduction on 5.29x in memory footprint was achieved compared to FP16. Note that in the LLaMA-3.2 series, it is necessary to separate the embedding table and head modules to satisfy the invariance arising from their tying. Furthermore, as the size of the embedding table has increased compared to previous models, the compression ratio has decreased accordingly.
5.3 Ablation Studies
#Bits Rotation LWC LDP PPL
2-4-4 17.40
8.93
10.59
8.31
Table 5: Ablation study on the impact of each component of our proposed RCP on performance for LLaMA-2 7B.
Impact of RCP Components
As shown in Table 5, we conducted an ablation study to analyze the impact of removing each component of RCP on model performance. In 4-bit activation quantization, addressing the outliers in activations was crucial, and this was effectively resolved using rotation, which led to the largest performance gain compared to the baseline. This demonstrates that rotation is a viable solution when quantizing activations to low bit-width.
However, we found that the narrow weight distribution caused by rotation hindered successful training of LWC. Specifically, when examining the training process with rotation applied during LWC training, the training loss curve exhibited instability. The combination of low bit-width quantization challenges and the difficulty in finding an optimal LWC required training stabilization, which was achieved by LDP. LDP reduced PPL from 10.59 to 8.31, demonstrating a clear advantage.
Impact of Rotation Configuration
W2A4KV4 PPL
RCP 8.31
-$R_{3}$ 8.48
-[$R_{2}$,$R_{3}$] 8.83
-[$R_{3}$,$R_{4}$] 12.24
-[$R_{2}$,$R_{3}$,$R_{4}$] 12.76
-[$R_{1}$,$R_{2}$,$R_{3}$,$R_{4}$] 25.05
Table 6: Ablation study on the impact of rotation configuration for LLaMA-2 7B.
Since the rotation requires additional processes before and after inference, we investigated the performance trend by incrementally adding rotation matrices ($R_{1}$,$R_{2}$,$R_{3}$,$R_{4}$) to different components to find an appropriate balance between accuracy and overhead. The results are presented in Table 6. The table demonstrates that the impact of the rotation was most significant with $R_{1}$ and $R_{4}$. Especially, $R_{1}$, which applies rotation matrix to the input weight and input activation of all modules thereby having the largest impact on quantization performance. Additionally, our analysis revealed that in LLaMA-2 7B, the input to the down projection layer (of the MLP) exhibited a significant number of outliers, which was effectively addressed through $R_{4}$ online rotation to activation.
6 Conclusion
RCP enables weights to be quantized to extreme low-bit precision through learnable non-uniform quantization while harmonizing with rotation to optimize both activations and KV-cache to 4-bit. RCP has achieved the first W2A4KV4 configuration and implemented optimized kernels for inference, facilitating LLM serving even in resource-constrained environments.
Limitations
Although our proposed RCP first enables challenging W2A4KV4 quantization of commonly used LLM models, we report key limitations of our work.
First, the online rotation operators ($R_{2}$ through $R_{4}$) inevitably introduce additional latency for training and evaluation. Custom CUDA kernels or FlashAttention3 Shah et al. (2024) can minimize such speed-down, however, it might not be a viable option for many edge application scenarios where no hardware support for fast Hadamard transform is available.
Second, RCP requires heavier hyperparameter tuning than BitDistiller since rotation tends to make the model weights more sensitive to the choice of learning rate. This can be prohibitive when a user is under a strict budget limit.
In future work, we could explore applying an optimized rotation matrix that achieves comparable performance to Cayley-optimized rotation matrices used in SpinQuant Liu et al. (2024b) while maintaining similar computational costs to the Random Hadamard rotation matrices employed in QuaRot Ashkboos et al. (2024b).
Appendix A Appendix
A.1 Related Works
PTQ and QAT
GPTQ Frantar et al. (2022) introduced an accurate post-training quantization (PTQ) method based on approximate second-order information that enables weight-only quantization down to 3-4 bits through block-wise reconstruction. SmoothQuant Xiao et al. (2023) proposed smoothing activation outliers by offline migrating quantization difficulty from activations to weights through equivalent transformation, enabling accurate 8-bit weight-activation quantization. AWQ Lin et al. (2024a) built upon SmoothQuant’s equivalent transformation concept but introduced activation-aware channel-wise scaling to protect salient weights during weight-only quantization. OmniQuant Shao et al. (2024) enhanced quantization by introducing learnable weight clipping and equivalent transformation parameters that are jointly optimized through block-wise reconstruction.
LLM-QAT Liu et al. (2024a) was the first to explore quantization-aware training (QAT) for LLMs using data-free knowledge distillation from the full-precision model to guide low-bit quantization. BitDistiller Du et al. (2024) improved upon LLM-QAT by introducing a self-distillation framework with confidence-aware KL divergence to enable sub-4-bit quantization while maintaining efficiency. EfficientQAT Chen et al. (2024b) made QAT more practical by introducing block-wise training of all parameters followed by end-to-end training of quantization parameters.
Rotation
QuaRot Ashkboos et al. (2024b) introduced a rotation-based approach using Hadamard transformations to eliminate outliers in activations and KV-cache, enabling end-to-end 4-bit quantization including weights, activations and KV-cache. SpinQuant Liu et al. (2024b) enhanced this rotation-based approach by learning optimal rotation matrices instead of using random ones.
Non-uniform Quantization
PACT Choi et al. (2018) introduced a learnable clipping parameter for activation quantization during training to help preserve model accuracy. SqueezeLLM Kim et al. (2024) took a different direction by focusing on identifying and extracting outlier values into a sparse format while quantizing the remaining dense values. NU2U Liu et al. (2022) proposed learning flexible non-uniform input thresholds while maintaining uniform output levels to balance quantization accuracy with hardware efficiency.
Serving Optimization
Atom Zhao et al. (2024) first introduced W4A4 quantization for LLM serving but faced performance challenges from dequantization overhead. QServe Lin et al. (2024b) addressed the challenges by introducing W4A8KV4 quantization with progressive group quantization FLUTE Guo et al. (2024) focused on developing efficient GPU kernels for flexible lookup table-based quantization methods that can support arbitrary bit widths including 3-bit and 4-bit quantization.
A.2 Reasoning Task Example: HumanEval
We evaluate the capability of the WizardCoder 7B model to generate solutions for coding problems. The results are presented in Fig. 8. The orange box in Fig. 8 represent the model output after applying rotation and quantizing the weights to W2A4KV4 using a uniform asymmetric quantizer. Under uniform quantization, it is evident that the model fails to perform logical generation tasks even applying rotation; it merely produces the structural template of code without generating functionality correct code. In contrast, the green box shows the results when the weights are quantized to W2A4KV4 using LDP. Unlike the uniform quantizer, the LDP approach yields code that not only adheres faithfully to the given instructions and generates a functionality correct algorithm, but also provides detailed explanatory comments. While perplexity on standard language modeling tasks did not reveal significant differences between the two cases, these findings suggest that LDP plays a crucial role in enabling logical reasoning tasks under extreme low-bit quantization.
A.3 Implementation Details
All model parameters are in BF16 format throughout training and evaluation since we observe overflow in the hidden activation of the last two FFNs on several models set to FP16.
In existing rotation-based PTQ methods Ashkboos et al. (2024b); Liu et al. (2024b), rotations are done in FP32 to avoid precision issues. However, this leads to computational overhead due to a large number of typecasting. When fusing rotations to model weights, they are temporarily promoted to FP32, multiplied by an appropriate rotation matrix, and then demoted back to their original precision. For online rotations ($R_{2}$, $R_{3}$, and $R_{4}$), all tensors are processed in BF16.
A.4 More Ablation Studies
Factorized Rotation
#Bits Factorized Batch Epoch PPL
W2 8 8 7.6
1 64 12.5
Table 7: Comparison of factorized configurations.
In our algorithm, rotation serves as a pre-conditioning tool for reducing outliers in activation and KV-cache. All rotations except the matrices that should be applied online ($R_{3}$ and $R_{4}$) are fused into the corresponding model weight at the beginning of the QAT process. This means their orthogonality is not guaranteed during backpropagation steps with AdamW optimizer.
We investigate the impact of preserving the orthogonality of the rotations by modifying the LLaMA-2 model implementation to apply all rotation operators online while freezing the rotation matrices. Table 7 presents the results. Applying factorized rotation prevents the fusion of the rotation matrix into the weight tensor, resulting in an increase in the number of intermediate tensors (rotation matrix and intermediate activation), which significantly raises VRAM requirements. For instance, applying only $R_{1}$ needs to reduce the training batch size from 8 to 1. Under the condition of maintaining an equal total number of tokens processed by the model, we compared the performance of W2A16KV16 with only $R_{1}$ applied. The perplexity of BitDistiller with $R_{1}$ fused was 7.6, whereas applying QAT with factorized rotation resulted in a PPL of 12.5. This indicates that performing weight updates through QAT while preserving $R1$ orthogonality hinders QAT optimization. This is because the factorization constrains the weight updates to a restricted space defined by the factorized condition, requiring the backpropagation process to maintain within this space. This limitation reduces the flexibility of optimization, making it challenging to efficiently adjust the weights. Consequently, this leads to suboptimal training dynamics and ultimately results in degraded model performance. Furthermore, extending factorization to $R_{2}$ and $R_{4}$ would lead to an even greater increase in VRAM usage. In contrast, training fused weight effectively alters only the distribution and is analogous to standard LLM training, which is well-known to perform effectively. In summary, given that resource consumption increases while performance degrades, we have decided not to explicitly preserve orthogonality and instead allow the algorithm to handle this aspect.
Layerwise vs. End-to-end QAT
Recent work introduced layerwise QAT Chen et al. (2024b), which updates one layer at a time while freezing others, allowing training on a single GPU. We extended this approach by applying rotation but observed significant performance degradation. The main issue stemmed from fuse rotation matrices in the weights; layerwise updates disrupted orthogonality, preventing the activation space from restoring its original space, leading to cumulative errors and reduced accuracy. In contrast, end-to-end methods like BitDistiller naturally mitigate this issue during updates. While factorized rotation could help, its high GPU memory requirements for holding rotation matrices and intermediate tensors on GPU memory offsets the advantage. Despite these challenges, exploring single GPU training using rotation matrix remains a promising direction for future work.
A.5 GEMM Kernel Design for Non-uniform W2A4 Quantization
[…]
A.6 Details and More Results on GEMV
(a) Dequantization process of two INT4 activations packed in INT8.
(b) Dequantization process of 4 UINT2 weights packed in UINT8.
Figure 7: Online dequantization of INT4 activations and UINT2 weights.
Online Dequantization and Vectorization
Fig. 7 illustrates how the activations and weights are dequantized in our GEMV kernel. For activations, there are two INT4 elements ($\text{X}_{hi},\text{X}_{low}$) in a packed INT8 $\text{X}_{q}$. For $\text{X}_{hi}$, $\text{X}_{q}$ is copied to an INT8 register, and the register is right-shifted by 4 bits with sign-filling. For $\text{X}_{low}$, $\text{X}_{q}$ is also copied to an INT8 register, which is left-shifted by 4 bits first to put the sign bit of $\text{X}_{low}$ to the MSB and then right-shifted by 4 bits with sign filling. This process is shown in Fig. 7(a).
For weights, there are four UINT2 elements ($\text{W}_{q0},\text{W}_{q1},\text{W}_{q2},\text{W}_{q3}$) in a packed UINT8 $\text{W}_{q}$. $\text{W}_{q}$ is copied to 4 UINT8 registers (for each UINT2 element) that are used as indices to look up the LUT $\hat{\textbf{W}}$. For $\text{W}_{q0}$, the register is right-shifted by 6 bits. For $\text{W}_{q1}$, the register is right-shifted by 4 bits, and a logical AND operation with a bit mask 0x03 is applied to select only two LSBs. For $\text{W}_{q2}$, the register is right-shifted by 2 bits and also performs logical AND with a bit mask 0x03. For $\text{W}_{q3}$, the register only does a logical AND with a bit mask 0x03.
The unit dequantization operations can be vectorized to increase memory throughput so that each thread writes 16B of data to shared memory. For activations, 4 $\text{X}_{q}$s are loaded from global memory at once by type casting via reinterpret_cast<char4 *>, which produces 8 FP16 dequantized activations to be written in sX. The dequantization is performed the same on each $\text{X}_{q}$ in a char4 struct. For weights, 2 $\text{W}_{q}$s are loaded from memory via reinterpret_cast<uint16_t *>. Unlike the activation case, the right-shift and logical AND operation can be naturally iterated 8 times to generate 8 FP16 dequantized weights that are directly multiplied to the corresponding activation from sX.
Shared Epilogue
As mentioned in Section 4.4, a shared output can be necessary due to our chunking strategy. For example, if BH is 4, then two warps will compute one output element to process a weight chunk of size BH/2 × C/4, and after warp-level sum reduction, the reduced values from the two warps must be summed once again. To implement this, we allocate a shared output buffer sO with twice the number of warps.
After the inner product stage for the first weight chunk, each thread in a block will have an FP32 accumulator with a shape of (4, 32). Applying the warp-shuffle primitive __shfl_xor_sync 5 times allows us to sum all accumulations to the first thread of each warp without any global nor shared memory access, producing 4 FP32 values to be cast to FP16 and stored in $\text{s}\textbf{O}[0:4]$. The first and the last two values are summed up as the first and the second output elements, respectively. Repeating the same process on the second weight chunk will produce the next 4 FP32 values for $\text{s}\textbf{O}[4:8]$ to compute the third and the fourth output elements accordingly.
Latency Benchmark
Our GEMV kernel is fully written in CUDA 12.1 and compiled for Nvidia A100 SXM 40GB model. We build our benchmarking framework upon QuaRot’s Ashkboos et al. (2024b) implementation that provides proper PyTorch bindings and a basic activation quantizer that combines a max reduction function written in PyTorch and a symmetric INT quantizer with INT4 sub-byte data handler from CUTLASS [^4] https://github.com/NVIDIA/cutlass.
[…]
Model #Bits (W-A-KV) Configuration PIQA Hella. Wino. ARC-c Avg.
Method Rotation LDP
1-7B 16-16-16 79.80 76.10 70.10 47.60 68.4
2-4-16 BitDistiller 61.53 35.98 49.25 23.46 43.56
Ours 70.67 45.86 62.03 30.54 52.28
Ours 70.62 46.41 61.48 31.32 52.46
2-4-4 BitDistiller 63.38 34.32 50.82 23.80 43.08
Ours 71.10 45.91 59.82 32.00 52.21
Ours 72.36 45.91 58.64 32.25 52.29
3-4-16 BitDistiller 73.34 50.94 63.61 34.81 55.68
Ours 76.71 53.96 68.19 35.23 58.52
Ours 77.20 53.11 68.43 38.82 59.39
3-4-4 BitDistiller 73.06 50.78 65.03 35.32 56.05
Ours 76.98 53.12 66.77 37.03 58.48
Ours 75.46 53.06 67.88 37.80 58.55
2-7B 16-16-16 77.86 57.14 68.35 43.34 61.67
2-4-16 BitDistiller 62.95 37.33 50.20 22.95 43.36
Ours 70.13 45.02 60.77 30.03 51.49
Ours 69.48 45.22 59.75 29.95 51.10
2-4-4 BitDistiller 62.70 37.18 53.91 25.93 44.93
Ours 69.53 45.67 59.35 29.86 51.10
Ours 69.91 44.58 59.70 30.69 51.22
3-4-16 BitDistiller 74.42 51.36 62.66 36.17 56.15
Ours 76.06 54.26 66.45 40.35 59.28
Ours 76.65 54.25 67.80 40.35 59.74
3-4-4 BitDistiller 72.41 50.51 63.29 35.83 55.51
Ours 76.55 53.55 65.90 39.33 58.83
Ours 76.71 53.88 65.43 41.04 59.27
3-8B 16-16-16 80.70 79.60 73.70 57.70 72.93
2-4-16 BitDistiller 57.23 29.96 49.48 21.16 39.46
Ours 69.96 44.30 59.43 28.66 50.59
Ours 69.16 44.67 59.91 29.69 50.86
2-4-4 BitDistiller 56.42 29.57 52.09 20.90 39.75
Ours 69.15 43.62 57.85 28.58 49.80
Ours 69.97 44.32 59.51 27.82 50.41
3-4-16 BitDistiller 72.47 49.72 62.43 36.94 55.39
Ours 77.25 55.18 68.90 42.91 61.06
Ours 77.64 55.21 69.93 43.34 61.53
3-4-4 BitDistiller 73.32 49.97 64.87 37.45 56.35
Ours 75.35 53.95 67.64 41.80 59.69
Ours 76.16 54.35 71.19 42.75 61.11
Table 8: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA models with parameter sizes of at least 7B.
Model #Bits (W-A-KV) Configuration PIQA Hella. Wino. ARC-c Avg.
Method Rotation LDP
1B 16-16-16 75.30 60.70 60.90 38.70 58.90
2-4-16 BitDistiller 51.95 27.41 48.46 19.45 36.82
Ours 61.15 30.66 50.67 21.84 41.08
Ours 61.42 31.55 51.78 20.65 41.08
2-4-4 BitDistiller 55.33 26.62 48.46 19.79 37.55
Ours 61.75 30.05 51.22 20.05 40.77
Ours 60.71 31.54 53.51 21.42 41.80
3-4-16 BitDistiller 53.53 28.35 48.61 19.62 37.53
Ours 69.53 40.31 55.40 26.27 47.88
Ours 69.64 40.57 56.12 26.37 48.18
3-4-4 BitDistiller 54.18 28.26 50.90 21.67 38.75
Ours 68.98 37.80 55.40 26.36 47.14
Ours 68.12 39.30 56.12 26.11 47.41
3B 16-16-16 76.00 71.00 66.60 47.60 65.30
2-4-16 BitDistiller 54.02 26.80 52.48 18.25 37.89
Ours 65.99 36.51 52.48 26.19 45.29
Ours 65.43 37.35 54.70 25.43 45.71
2-4-4 BitDistiller 51.84 26.70 51.38 19.11 37.26
Ours 64.30 36.26 51.38 25.08 44.26
Ours 65.45 36.66 53.75 26.37 45.56
3-4-16 BitDistiller 52.72 26.66 50.43 19.45 37.32
Ours 74.04 49.56 63.22 35.83 55.66
Ours 73.77 49.52 62.65 37.54 55.87
3-4-4 BitDistiller 53.91 26.82 48.03 20.30 37.27
Ours 74.31 49.19 60.06 36.77 55.08
Ours 73.18 48.87 62.43 36.01 55.12
Table 9: Complete comparison of accuracy on Zero-shot Common Sense Reasoning tasks on LLaMA-3.2 models with parameter sizes of at most 3B.
Layer Size FP16 RCP RCP+FP16Had RCP+FP32Had QuaRot QuaRot+FP16Had QuaRot+FP32Had
(2048, 8192) 0.054 0.036 0.110 0.146 0.073 0.155 0.186
(3072, 8192) 0.054 0.035 0.169 0.198 0.074 0.212 0.237
(4096, 11008) 0.077 0.048 0.120 0.148 0.088 0.157 0.186
(4096, 14336) 0.110 0.059 0.121 0.149 0.079 0.157 0.183
Table 10: GEMV latency for the down projection is measured except activation quantization overhead. The layer size is composed as (input channel, output channel). All latency numbers are in milliseconds.
Layer Size RCP RCP+FP16Had RCP+FP32Had QuaRot QuaRot+FP16Had QuaRot+FP32Had
(2048, 2048) 0.131 0.214 0.248 0.170 0.248 0.276
(3072, 3072) 0.131 0.265 0.295 0.168 0.304 0.331
(4096, 4096) 0.133 0.221 0.250 0.174 0.250 0.282
Table 11: GEMV latency for the value projection is measured including activation quantization overhead. The layer size is composed as (input channel, output channel). All latency numbers are in milliseconds.
Layer Size RCP RCP+FP16Had RCP+FP32Had QuaRot QuaRot+FP16Had QuaRot+FP32Had
(2048, 8192) 0.143 0.218 0.240 0.186 0.261 0.289
(3072, 8192) 0.140 0.271 0.294 0.177 0.318 0.340
(4096, 11008) 0.143 0.223 0.250 0.177 0.264 0.288
(4096, 14336) 0.142 0.226 0.247 0.177 0.259 0.285
Table 12: GEMV latency for the down projection is measured including activation quantization overhead. The layer size is composed as (input channel, output channel). All latency numbers are in milliseconds.
Figure 8: A reasoning task example from HumanEval  Chen et al. (2021) benchmark, conducted by two differently quantized WizardCoder-7B  Luo et al. (2023) models. The results in the orange box is from state-of-the-art QAT method BitDistiller Du et al. (2024) with applying rotation. In the green box, our proposed RCP is applied. Both methods employ exactly the same 4-bit quantization setting for activation and KV-cache.
A.7 Information About Use of AI Assistants
AI assistance was strictly limited to linguistic perspectives, such as grammar and spell checking, and finding synonyms.